OpenStudy (anonymous):

Together, teammates Pedro and Ricky got 2672 base hits last season. Pedro had 280 more hits than Ricky. How many hits did each player have?

OpenStudy (anonymous):

May I help you??

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So first we must create an equation Generally we know that Pedro and Ricky combined have a total of 2672 baseball hits. (To make it easier let's make P - Pedro and R- Ricky) So our general equation is P+R = 2672 Do you agree??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok So with the info given we know that P has 280 hits more than R so do you have an idea of an equation we can create?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok that's ok. To determine what P's total is we can use the equation P = R + 280

OpenStudy (anonymous):

Are you understanding?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so we plug in our equation for P in our general equation so can you tell me what that is?

OpenStudy (anonymous):

p+r=280

OpenStudy (anonymous):

Yes now replace P with the equation we just found

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can I see?? :P :S

OpenStudy (anonymous):

Do you need help?

OpenStudy (anonymous):

no am calculating thank you so much

OpenStudy (anonymous):

Ok So do you still need my help or are you super on your own??

OpenStudy (anonymous):

no thank you so much for your help i get it

OpenStudy (anonymous):

Yay Congratulations!

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

Ok What is it?